# Strand, London WC2R 2LS, UK

### v1

@@ Appendix G Lorentzian Linearized Dynamics in Einstein-Maxwell Gravity
Here we briefly discuss the behavior of linearized modes subject to conformal boundary conditions in Einstein-Maxwell gravity, about backgrounds with planar and hyperbolic horizons. [^33] The case of uncharged spherical horizons has been analyzed previously in refs. Anninos:2023epi ; Anninos:2024wpy ; Anninos:2024xhc ; Liu:2024ymn . The addition of charge does not affect their conclusions, at least for the $l=0$ modes. Overall, these modes exhibit better stability properties than the corresponding modes of spherical horizons.
@@
and our boundary placed at $r=e^{{\boldsymbol{\omega}}(u)}\mathfrak{r}$. The CBC (3) read: [^34] Note that here we are using a slightly different metric ansatz for the black brane compared to the one in eq. (80), to make contact with the known literature, e.g., ref. Anninos:2024xhc .
@@
Requiring $\mathcal{K}$ to be constant yields a differential equation for the Weyl factor ${\boldsymbol{\omega}}(u)$:
@@
which in turn admits solutions of the form:
@@
where $\mathcal{K}$ here corresponds to the extrinsic curvature of $\partial\mathcal{M}$ for the conformal representative metric with ${\boldsymbol{\omega}}(u)=0$:
@@
The dynamics of the conformal factor are governed by the following differential equation:
@@
which upon linearization reduces to: [^35] Note that this equation is identical to the one for spherical horizons Liu:2024ymn ; Galante:2025tnt and therefore the only information about the (spherical or hyperbolic) nature of the horizon comes through $f(r)$.
@@
and thus admits solutions of the form:

### v2

@@ Appendix G Lorentzian Linearized Dynamics in Einstein-Maxwell Gravity
Here we briefly discuss the behavior of linearized modes subject to conformal boundary conditions in Einstein-Maxwell gravity, about backgrounds with planar and hyperbolic horizons. [^34] The case of uncharged spherical horizons has been analyzed previously in refs. Anninos et al. (2023a, 2024, 2025a); Liu et al. (2024). The addition of charge does not affect their conclusions, at least for the $l=0$ modes. Overall, these modes exhibit better stability properties than the corresponding modes of spherical horizons.
@@
and our boundary placed at $r=e^{{\boldsymbol{\omega}}(u)}\mathfrak{r}$. The CBC (3) reads [^35] Note that here we are using a slightly different metric ansatz for the black brane compared to the one in eq. (80), to make contact with the known literature, e.g., ref. Anninos et al. (2025a).
@@
Requiring $\mathcal{K}$ to be constant yields a differential equation for the Weyl factor ${\boldsymbol{\omega}}(u)$,
@@
which in turn admits solutions of the form,
@@
where $\mathcal{K}$ here corresponds to the extrinsic curvature of $\partial\mathcal{M}$ for the conformal representative metric with ${\boldsymbol{\omega}}(u)=0$,
@@
The dynamics of the conformal factor are governed by the following differential equation,
@@
which upon linearization reduces to [^36] Note that this equation is identical to the one for spherical horizons Liu et al. (2024); Galante et al. (2025) and therefore the only information about the (spherical or hyperbolic) nature of the horizon comes through $f(r)$.
@@
and thus admits solutions of the form